Circular No. 7270

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

CASSIOPEIA A G. G. Pavlov, Pennsylvania State University; and V. E. Zavlin, Max-

Planck-Institut für Extraterrestrische Physik, report: “The analyses of two observations of Cas A taken with the Einstein HRI on 1979 Feb. 8 and 1981 Jan. 22 show a pointlike source at α = 23^{h}23^{m}27^{s}.85 ± 0^{s}.05, δ = +58^{o}48′43′′.8 ± 0′′.1, (equinox 2000.0). This appears to be the same object reported by Tananbaum et al. from the Chandra ACIS observation (IAUC 7246) and by Aschenbach from the ROSAT HRI observation (IAUC 7249). The apparent separations between the Einstein and Chandra positions of the source, 1′′.8 and 2′′.6, respectively, are comparable with the position uncertainties, which indicates that the source has not moved appreciably during the 20-yr span of observation. The Einstein HRI background-corrected countrates are 0.0009 ± 0.0002 and 0.0011 ± 0.0003 counts/s (exposure times 42 700 and 25 600 s), respectively. The Einstein HRI, ROSAT HRI, and Chandra ACIS countrates are consistent with the assumption that the source spectral flux was the same at each of the four epochs. If the source radiates as a blackbody, its size is too small (and the temperature too high) to interpret the radiation as emitted from the whole surface of an isolated, cooling neutron star.”

COMET C/1999 J3 (LINEAR)

Ephemeris extension to IAUC 7252 (elements MPC 35814):

1999TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Oct. 9 7^{h}53.^{m}65 + 6^{o}06.^{′}6 0.576 1.030 76.^{o}5 70.^{o}6 7.5
11 7 49.36 + 0 24.4 0.550 1.041 78.7 70.1 7.5
13 7 44.65 − 5 47.5 0.530 1.054 80.9 69.3 7.5
15 7 39.42 −12 23.7 0.516 1.067 83.1 68.2 7.5
17 7 33.58 −19 15.6 0.509 1.081 85.2 66.8 7.5
19 7 27.01 −26 11.8 0.509 1.096 87.1 65.2 7.6
21 7 19.58 −32 59.9 0.517 1.111 88.7 63.6 7.8
23 7 11.12 −39 28.7 0.532 1.128 90.0 61.9 7.9
25 7 01.44 −45 29.2 0.553 1.145 90.9 60.3 8.1
27 6 50.31 −50 55.9 0.579 1.163 91.4 58.7 8.3
29 6 37.49 −55 46.1 0.611 1.181 91.7 57.2 8.5

Visual m_{1} estimates: Oct. 2.07 UT, 7.9 (B. H. Granslo, Fjellhamar, Norway, 0.07-m refractor); 5.78, 7.5 (Y. Nagai, Yamanashi, Japan, 11×80 binoculars); 7.35, 7.7 (J. Bortle, Stormville, NY, 20×80 binoculars).

1999 October 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green